OpenStudy (anonymous):

At what point on the paraboloid y = x^2+z^2 is the tangent plane parallel to the plane x+2y+3z=1?

OpenStudy (turingtest):

find the gradient of the surface

OpenStudy (anonymous):

<2,-1,2>

OpenStudy (turingtest):

grad f=<2x, -1, 2z>

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

ok, now since we can't plug in anything that will change y, let's multiply the gradient by a scalar to get the j-component equal to 2 like it is for the plane

OpenStudy (anonymous):

multiply by -2

OpenStudy (turingtest):

right, and so we get?

OpenStudy (anonymous):

grad f=<-4x,2y, -4z>

OpenStudy (turingtest):

grad f=<-4x,2,-4z> that y can't just appear from nowhere

OpenStudy (turingtest):

so what value of x makes the i-components equal for the plane and grad f ? what value of z makes the k-components equal?

OpenStudy (anonymous):

x=1/2and z= 3/2 ???

OpenStudy (turingtest):

from the equation as we have it we need x=-1/4 and z=-3/4

OpenStudy (anonymous):

ok.

OpenStudy (turingtest):

that's all then... boring problem

OpenStudy (anonymous):

why qns with sphere have 2 sets of points?

OpenStudy (turingtest):

hm... seems strange that y does not matter. Maybe I messed this one up :P

OpenStudy (turingtest):

sorry? what do you mean?

OpenStudy (anonymous):

for qns with sphere has 2 sets of points as ans why?

OpenStudy (turingtest):

what do you mean it has 2 sets of points?

OpenStudy (anonymous):

the ans fot this qn is (-1/4, -2,-3/4)

OpenStudy (turingtest):

the way we did it it looks like y does not matter, so (-1/4,y,-3/4) seems strange that y does not matter though, maybe I made a mistake... not sure

OpenStudy (anonymous):

so the point for y is 0 or -1?

OpenStudy (turingtest):

neither, the result says y can be anything, so just write y=y

OpenStudy (amistre64):

isnt y a function if x and z to begin with? y = x^2 + z^2; therefore y=1/4